# Towards Better Visualizing the Decision Basis of Networks via Unfold and Conquer Attribution Guidance

### v1

@@ Introduction
In terms of explainability, it is crucial to provide the intuitive and convincing visualizations from a human-view. To demonstrate how our method improves the quality of generated explanations, we compare the results with the representative explanation methods in Figure  4. The first and second rows represent the results of original and our methods, respectively. Red and blue colors denote the positive and negative relevance scores, respectively. While existing methods have the advantage of depicting the high relevant regions to the output predictions, they still lack providing fine-grained descriptions with the limited granularity of representing areas of contribution.
As the manner of post-hoc framework, UCAG has the strong advantages of fine-grained visualization, easily applicable and agnostic to model variations. In visualization, it improves the quality of existing maps by revising the incorrect negative or positive relevance scores in the (ir)relevant area without losing salient regions, resulting in increasing the density of the explanation map. Figure 5 shows the improvement when UCAG applied to various models: ResNet, DenseNet, and InceptionNet. These visual enhancements also lead to improved quantitative performance. We report the additional visual explanations and evaluations on various methods and models, including GradCAM, FullGrad, GradCAM++, WGradCAM, RAP, RSP, AGF and SGLRP, transformer interpretability .
@@
Table 1: The AUC scores regarding the deletion (lower is better)/insertion (higher is better) games on ImageNet dataset. Mark* represents the performance of applying our methods.
@@
Methods
VGG16
ResNet50
VGG16
ResNet50
Center
69.6/42.4
69.6/42.4
27.8/19.5
27.8/19.5
Gradient
76.3/56.9
72.3/56.8
37.7/31.4
35.0/29.4
DeConv
67.5/44.2
68.6/44.7
30.7/23.0
30.0/21.9
Guid
75.9/53.0
77.2/59.4
39.1/31.4
42.1/35.3
MWP
77.1/56.6
84.4/70.8
39.8/32.8
49.6/43.9
cMWP
79.9/66.5
90.7/82.1
49.7/44.3
58.5/53.6
RISE
86.9/75.1
86.4/78.8
50.8/45.3
54.7/50.0
GradCAM
86.6/74.0
90.4/82.3
54.2/49.0
57.3/52.3
Extremal
88.0/76.1
88.9/78.7
51.5/45.9
56.5/51.5
NormGrad
81.9/64.8
84.6/72.2
CAMERAS
86.2/76.2
94.2/88.8
55.4/50.7
69.6/66.4
@@ Methods
Table 2: The performance of the pointing game among various methods. Our method (apply to the GradCAM) represents the sizeable increment of performance compared to the other method.
@@
The deletion and insertion game measures whether the explanation map correctly highlights areas contributing to the decision-making process. Deletion measures the model’s response by repeatedly deleting the high-priority pixels with respect to the output. When deleted pixels contribute significantly to predictions, model confidence and deletion scores decrease dramatically. In contrast, insertion game measures how the confidence score would be increased. we set the baseline image blurred with kernel size: $51$ and $\sigma$: $50$ for the insertion game and images with zero pixels for the deletion game.
@@
Table 3: The performance of the energy–based pointing game (all/diff) and among various methods and datasets.
@@
VGG mAP 0.789/0.783 0.700/0.682 0.534/0.559 0.547/0.546 0.735/0.743 0.689/0.709 0.740/0.745
Pixel Acc. 0.742/0.737 0.750/0.730 0.547/0.556 0.786/0.789 0.725/0.742 0.737/0.756 0.779/0.789
ResNet mAP 0.748/0.774 0.717/0.722 0.671/0.694 0.671/0.694 0.711/0.761 0.713/0.760 0.758/0.782
Pixel Acc. 0.798/0.821 0.770/0.772 0.716/0.736 0.716/0.722 0.763/0.809 0.766/0.807 0.804/0.824
Table 4: Performance of mAP and outside–inside ratio over ImageNet segmentation dataset (Original / Ours). Attribution methods that have the class-discriminative ability are compared. Low Out/In ratio denotes attributions are intensively distributed in the semantic mask.
@@
Evaluating localization performance is employed to assess the quality of the explanation, assuming that the highlighted pixels contains the main object of input image. (zhang2018top) introduced Pointing Game, widely used in many related researches, that computes whether the maximum point of the saliency map is contained in the label. We report the result of ours (GradCAM with UCAG) in Table 2 and it outperforms the state-of-art explanatory approaches.
ScoreCAM (wang2020score) raised the issue that pointing game does not provide a detailed measure of localization performance since it only considers the maximum saliency point inside the explanation. As an alternative, ScoreCAM introduces an energy–based pointing game to evaluate the quality with a more detailed measurement. The energy–based pointing game is based on the summation of the regions where saliency overlaps the inside label. We also report the comparisons of the energy–based pointing game in Table 3 with the results of various CAM approaches. The increased scores indicate that UCAG consistently enhances the ability of various CAM approaches regardless of the type of network and method.
Evaluating the weakly supervised segmentation is closely related to interpretability in terms of finding the salient objects in input image. In recent works (nam2020relative; nam2021interpreting; gur2021visualization), mean average precision (mAP) is utilized to assess the quality of interpretability. We follow the experiential settings with ImageNet segmentation dataset (guillaumin2014imagenet). Table 4 shows the comparisons of existing methods and ours. In most cases, our method improves the quality of interpretations in both VGG and ResNet models.
Model
GCAM CAMERAS Ours
Positive map density $(p_{map}^{+}\uparrow{})$
ResNet50
2.33 3.20 3.89
DenseNet
2.35 3.23 3.45
Inceptionv3
2.18 3.15 3.83
Negative map density $(p_{map}^{-}\downarrow{})$
ResNet50
0.86 0.81 0.81
DenseNet
0.94 0.83 0.82
Inceptionv3
1.04 0.93 0.85
Table 5: Evaluated results of positive (higher is better) and negative (lower is better) map densities.
@@
Several previous works (jalwana2021cameras; poppi2021revisiting) pointed out that the drop and increase in confidence do not reflect the fine-grained representation ability of the explanation map. As an alternative, CAMERAS (jalwana2021cameras) introduces the positive and negative map densities metric, which concurrently evaluates the change in model confidence computed from the perturbed image with saliency map and density. positive and negative map density are defined as:
$\rho_{map}^{+}=\frac{\psi(f(M{\odot}X))\times(h\times w)}{\sum_{i}{\sum_{j}{M^{(i,j)}}}},$ (7)
$\rho_{map}^{-}=\frac{\psi(f(1-M{\odot}X))\times(h\times w)}{\sum_{i}{\sum_{j}{(1-M^{(i,j)})}}}.$ (8)
Since the metric provides the measures of the density as well as an increase in the prediction of the model, the precise quality of the saliency maps is quantitatively measured.
The results contained in table 5 show that UCAG enhances the density of saliency maps more than the multi-resolution fusion approach. This implies that our approach has the advantage of strong objectness with the concentrated scores in the saliency regions.
@@
In this section, we discuss the effectiveness of spatial scrutinization performed by UCAG. For comparison, performance with and without (only takes the scaled image as input) spatial scrutinization is measured in terms of causality, localization, and density. As shown in the figure, the insertion score remains the same as the deletion score decreases. In other words, spatial scrutinization derives the generation of saliency maps from preserving the areas contributing to the decision process. The density of the saliency map becomes higher than the naive image scaling approach while localizing the object is better than the image scaling approach. Hence, the saliency map generated by spatial scrutinization more precisely points to the actual causal pixels with increased density of the saliency map.
@@ Conclusion
This paper propose a novel post-hoc method to enhance the interpretability of network decision with the perspective of spatial scrutinization in input features. To fulfill this, we design a framework that encourages the CAM module to scrutinize images along spatial axes by providing partial pieces of evidence from a single ground image and aggregating the explanations generated from multiple perspectives. Qualitative and quantitative results show an remarkable performance improvement. Especially, the increased performance in causal and density matrices means that the explanation provides the causal pixels with a large density, which naturally improves localization performance. Therefore, it is reasonable to say that UCAG provides more clarified and clear explanations for the various explanation methods and models.

### v2

@@ Introduction
In terms of explainability, it is crucial to provide intuitive and convincing visualizations from a human view. To demonstrate how our method improves the quality of generated explanations, we compare the results with the representative explanation methods in Figure  4. The first and second rows represent the results of the original and our methods, respectively. Red and blue colors denote the positive and negative relevance scores, respectively. While existing methods have the advantage of depicting the highly relevant regions to the output predictions, they still lack providing fine-grained descriptions with the limited granularity of representing areas of contribution.
In the manner of the post-hoc framework, UCAG has the strong advantages of fine-grained visualization, easily applicable and agnostic to model variations. In visualization, it improves the quality of existing maps by revising the incorrect negative or positive relevance scores in the (ir)relevant area without losing salient regions, resulting in increasing the density of the explanation map. Figure 5 shows the improvement when UCAG is applied to various models: ResNet, DenseNet, and InceptionNet. These visual enhancements also lead to improved quantitative performance. We report the additional visual explanations and evaluations on various methods and models, including GradCAM, FullGrad, GradCAM++, WGradCAM, RAP, RSP, AGF, SGLRP, and transformer interpretability  (Chefer, Gur, and Wolf 2021).
@@
Table 1: The AUC scores regarding the deletion (lower is better)/insertion (higher is better) games on the ImageNet dataset. Mark * represents the performance of applying our methods.
@@
Methods VGG16 ResNet50 VGG16 ResNet50
Center 69.6/42.4 69.6/42.4 27.8/19.5 27.8/19.5
Gradient 76.3/56.9 72.3/56.8 37.7/31.4 35.0/29.4
DeConv 67.5/44.2 68.6/44.7 30.7/23.0 30.0/21.9
Guid 75.9/53.0 77.2/59.4 39.1/31.4 42.1/35.3
MWP 77.1/56.6 84.4/70.8 39.8/32.8 49.6/43.9
cMWP 79.9/66.5 90.7/82.1 49.7/44.3 58.5/53.6
RISE 86.9/75.1 86.4/78.8 50.8/45.3 54.7/50.0
GradCAM 86.6/74.0 90.4/82.3 54.2/49.0 57.3/52.3
Extremal 88.0/76.1 88.9/78.7 51.5/45.9 56.5/51.5
NormGrad 81.9/64.8 84.6/72.2 - -
CAMERAS 86.2/76.2 94.2/88.8 55.4/50.7 69.6/66.4
@@
Table 2: The performance of the pointing game among various methods. Our method (applied to the GradCAM) represents a sizeable increment in performance compared to the other method.
@@
The deletion and insertion game measures whether the explanation map correctly highlights areas contributing to the decision-making process. Deletion measures the model’s response by repeatedly deleting the high-priority pixels with respect to the output. When deleted pixels contribute significantly to predictions, model confidence and deletion scores decrease dramatically. In contrast, the insertion game measures how the confidence score would be increased. We set the baseline image blurred with kernel size: $51$ and $\sigma$: $50$ for the insertion game and images with zero pixels for the deletion game.
@@
Table 3: The performance of the energy-based pointing game represented as original / ours. ()(-)( - ) denotes the result in the difficult set of each dataset.
@@
VGG mAP 0.789/0.783 0.700/0.682 0.547/0.559 0.551/0.570 0.735/0.743 0.743/0.752 0.740/0.752
Pixel Acc. 0.742/0.737 0.750/0.730 0.534/0.556 0.543/0.587 0.786/0.789 0.793/0.796 0.779/0.794
ResNet mAP 0.748/0.774 0.717/0.722 0.671/0.694 0.588/0.617 0.711/0.761 0.719/0.766 0.758/0.783
Pixel Acc. 0.798/0.821 0.770/0.774 0.716/0.736 0.608/0.648 0.763/0.809 0.775/0.813 0.804/0.826
Table 4: Performance of mAP and outside–inside ratio over ImageNet segmentation dataset (Original / Ours). Propagation-based methods that have the class-discriminative ability are compared. A Low Out/In ratio denotes attributions are intensively distributed in the semantic mask.
@@
Evaluating localization performance is employed to assess the quality of the explanation, assuming that the highlighted pixels contain the main object of the input image. (Zhang et al. 2018) introduced the Pointing Game, widely used in many related researches, that computes whether the maximum point of the saliency map is contained in the label. We report our results (GradCAM with UCAG) in Table 2, and it outperforms the state-of-art explanatory approaches.
ScoreCAM (Wang et al. 2020) raised the issue that the pointing game does not provide a detailed measure of localization performance since it only considers the maximum saliency point inside the explanation. As an alternative, ScoreCAM introduces an energy-based pointing game to evaluate the quality with a more detailed measurement. The energy-based pointing game is based on the summation of the regions where saliency overlaps the inside label. We also report the comparisons of the energy-based pointing game in Table 3 with the results of various CAM approaches. The increased scores indicate that UCAG consistently enhances the ability of various CAM approaches regardless of the type of network and method.
Evaluating the weakly supervised segmentation is closely related to interpretability in terms of finding the salient objects in the input image. In recent works (Nam et al. 2020; Nam, Choi, and Lee 2021; Gur, Ali, and Wolf 2021), mean average precision (mAP) is utilized to assess the quality of interpretability. We follow the experiential settings with the ImageNet segmentation dataset (Guillaumin, Küttel, and Ferrari 2014). Table 4 shows the comparisons of existing methods and ours. In most cases, our method improves the quality of interpretations in both VGG and ResNet models.
Model GCAM CAMERAS Ours
Positive map density $(\mathbb{D}_{map}^{+}\uparrow{})$
ResNet50 2.33 3.20 3.89
DenseNet 2.35 3.23 3.45
Inceptionv3 2.18 3.15 3.83
Negative map density $(\mathbb{D}_{map}^{-}\downarrow{})$
ResNet50 0.86 0.81 0.81
DenseNet 0.94 0.83 0.82
Inceptionv3 1.04 0.93 0.85
Table 5: Evaluated results of positive (higher is better) and negative (lower is better) map density.
@@
Several previous works (Jalwana et al. 2021; Poppi et al. 2021) pointed out that only reporting the ratio of (de)increase in confidence score does not fully reflect the interpretability of explaining methods. As an alternative, CAMERAS (Jalwana et al. 2021) introduces the positive (negative) map density, which simultaneously evaluates the variations of confidence scores and saliency density during the perturbation. Metrics for positive (negative) map density are defined as $\mathbb{D}_{map}^{+}=f(I{\odot}\mathbb{M}^{k})/\sum_{i}{\sum_{j}{\mathbb{M}^{k}_{(i,j)}}}\times(h\times w)$ and $\mathbb{D}_{map}^{-}=f(I{\odot}1-\mathbb{M}^{k})/{\sum_{i}{\sum_{j}{(1-\mathbb{M}^{k}_{(i,j)})}}}\times(h\times w)$. Therefore, it is possible to measure the quality of the explanation in a more precise way.
The results in Table 5 show that UCAG enhances the density of saliency maps more than the existing multi-resolution fusion approach. This implies that our approach has the advantage of strong objectness with concentrated scores in the saliency regions.
@@
As an ablation study, the gap of effectiveness between upscaling and spatial scrutinization is described in Figure 2. UCAG provides a more precise explanation with the tendency to preserve the areas contributing to the decision-making process.
Here, we note that the computational cost is inevitably increased due to the manner of post-hoc framework. However, it is more efficient to find various spatial saliency with the scrutinization rather than the cost of calculating the exponentially increasing image resolution. In addition, we implement the overall frameworks batch-wisely, leading to enhancing computational efficiency.
@@ Conclusion
In this paper, we propose a post-hoc framework: UCAG, to enhance the interpretability of network decisions by guiding the explaining methods to carefully investigate the saliencies with spatial scrutinization and aggregation. We demonstrate the validity of UCAG from qualitative and quantitative results. With the advantage of being agnostic to the model and method, the increments of causality, localization, and density demonstrate the effectiveness of UCAG for producing a better visual representation ability of network decisions.
Acknowledgements
This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No.2022-0-00984, Development of Artificial Intelligence Technology for Personalized Plug-and-Play Explanation and Verification of Explanation & No. 2019-0-00079, Artificial Intelligence Graduate School Program, Korea University)